# Sard properties for polynomial maps in infinite dimension

### v2

@@ 3.3. Variations and their behaviour under polynomial maps
Exactly as in [YC04, Thm. 7.2], we cannot apply directly the area formula to $A$, since the dimension of $A$ can be larger than $i$. In [YC04, Thm. 7.2] this problem is solved using [YC04, Ex. 4.11]. However, the dependence on $n$ of the constants appearing there is not explicit. It could in be principle be made so, using the Tarski-Seidenberg theorem, but the dependence obtained in this way would be doubly exponential in $n$, which is not enough for our purposes. We overcome the obstacle using Theorem 13.
For sufficiently small $\varepsilon>0$ let $C_{\varepsilon}$ the semialgebraic set obtained by applying Theorem 13 to the set $A$ and the polynomial map $F=p_{Z}:\mathbb{R}^{n}\to\mathbb{R}^{i}$. We then proceed using the area formula to estimate $\mathcal{H}^{i}(\pi_{Z}\circ p(C_{\varepsilon}))$. (We will see at the end of the proof how $\mathcal{H}^{i}(\pi_{Z}\circ p(C_{\varepsilon}))\to\mathcal{H}^{i}(\pi_{Z}\circ p(A))$ as $\varepsilon\to 0$). By Item i of Theorem 13, $\dim(C_{\varepsilon})\leq i$. We can assume without loss of generality that $C_{\varepsilon}\subset\mathbb{R}^{n}$ is a smooth embedded submanifold with dimension equal to $i$, since the other strata give zero contribution to $\mathcal{H}^{i}$. Denoting with $\bar{p}_{Z}:C_{\varepsilon}\to\mathbb{R}^{i}$ the restriction, we obtain
@@
(84) $\leq\int_{C_{\varepsilon}}\sigma_{1}(D_{x}\bar{p}_{Z})\cdots\sigma_{i}(D_{x}\bar{p}_{Z})\mathcal{H}^{i}(dx).$ by definition of singular values
Note that, denoting with $j:C_{\varepsilon}\hookrightarrow\mathbb{R}^{n}$ the inclusion, we have
@@
(88) $\leq\sup_{x\in\mathcal{U}_{\varepsilon}(A)}\sigma_{1}(D_{x}p)\cdots\sigma_{i}(D_{x}p)\mathcal{H}^{i}(C_{\varepsilon})$ by Item ii of Theorem 13
(89) $\leq\left(\overline{\sigma}_{i}+\eta(\varepsilon)\right)\mathcal{H}^{i}(C_{\varepsilon}),$ by continuity of singular values
where $\overline{\sigma}_{i}$ is defined in (75), and $\eta(\varepsilon)\to 0$ as $\varepsilon\to 0$. Since $C_{\varepsilon}$ is a bounded semialgebraic set of dimension $i$, then $\mathcal{H}^{i}(C_{\varepsilon})=V_{i}(C_{\varepsilon})$ by Lemma 15. Hence
@@
(91) $\leq c(n,i)\beta_{\mathrm{dc}}^{n-i}\int_{G_{i}(\mathbb{R}^{n})}\left(\int_{Z}\mathbbm{1}_{\pi_{Z}(C_{\varepsilon})}(x)\,\mathcal{H}^{i}(dx)\right)\gamma_{i,n}\left(dZ\right)$ by Item iv of Theorem 13
(92) $\leq c(n,i)\beta_{\mathrm{dc}}^{n-i}\int_{G_{i}(\mathbb{R}^{n})}\left(\int_{Z}\mathbbm{1}_{\pi_{Z}(\mathcal{U}_{\varepsilon}(A))}(x)\,\mathcal{H}^{i}(dx)\right)\gamma_{i,n}\left(dZ\right)$ by Item ii of Theorem 13
@@
where $\beta_{\mathrm{dc}}=\beta_{\mathrm{dc}}(c,d,i)>1$ is the constant from Theorem 13, which (up to taking the maximum for $i\leq m$) depends only on $c,d,m$. Summing up, we have proved that for all $i=1,\dots,m$ it holds
@@
In order to take the limit in (95), we recall two properties of semialgebraic sets. The first one is the continuity of the Lebesgue measure in the Hausdorff topology in the semialgebraic category (see [YC04, Thm. 5.10]): if $\{S_{\varepsilon}\}_{\varepsilon\geq 0}\subset\mathbb{R}^{i}$ is a one-parameter family of bounded semialgebraic sets then
@@
The second property is the behaviour of variations for $\varepsilon$-neighbourhoods (see [YC04, Thm. 5.11]): if $A\subset\mathbb{R}^{n}$ is a bounded semialgebraic set, then for all $i=1,\dots,n$ it holds
@@
Therefore, using (96) (and recalling that by Item iii of Theorem 13 it holds $p_{Z}(C_{\varepsilon})\to p_{Z}(A)$ in the Hausdorff topology) and (97) we can pass to the limit in (95) and obtain
@@
To conclude, we estimate $V_{i}(A)$. Similarly as in (81), and recalling that $A\subset B_{\mathbb{R}^{n}}(r)$ we have
@@ 3.4. Variations of almost-critical values of polynomial maps
In this section we recall the notion of almost-critical points $C^{\Lambda}(p)$ of a polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$. The set $p(C^{\Lambda}(p))$ is the set of its almost-critical values. The main result of this section is Theorem 23, which provides a version of [YC04, Cor. 7.4] where the dependence of the parameters with respect to $n$ is explicit. This result is based on Theorem 18 and on the specific semialgebraic structure of almost-critical points of polynomial maps, that we prove in Proposition 22.
@@
where $E_{i,b_{i}+1}=\{(x,y)\in\mathbb{R}^{s}\times\mathbb{R}\mid y-t\leq 0\}$. In particular, this set admits a diagram that does not depend on $t$, since the degree of the polynomial $y-t$ is equal to $1$. As explained in Section 3.1.3, the projection of a semialgebraic set is semialgebraic and its diagram is controlled explicitly by the diagram of the original set only. In particular, the diagram of the projection (105) admits a description with a diagram $(s,I,J)$, that does not depend on $t$. ∎
@@
The functions $\lambda_{i}:{\operatorname{Sym}}_{m}^{+}\to\mathbb{R}$ and are continuous and semialgebraic.
@@
By Lemma 21, for every $k=1,\ldots,m$, the function $\lambda_{k}:{\operatorname{Sym}}_{m}^{+}\to\mathbb{R}$ is semialgebraic, hence from Lemma 20 applied to $A=\operatorname{Sym}_{m}^{+}$ and the semialgebraic functions $\lambda_{k}:A\to\mathbb{R}$, we deduce the following fact: for any $t\in\mathbb{R}$ we can write
@@
Now we state and prove the main result of this section (Theorem G in the Introduction).
@@
By Proposition 22 the set $A=C^{\Lambda}(p)\cap B_{\mathbb{R}^{n}}(r)$ is a closed and bounded semialgebraic set whose diagram is $D(A)=(n,c^{\prime},d^{\prime})$ where $c^{\prime}=c^{\prime}(d,m)$, $d^{\prime}=d^{\prime}(d,m)$ (which in particular do not depend on $n$, $\Lambda$). Furthermore, for the polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ it holds, by construction
@@
We can then apply Theorem 18, yielding (113) with $\beta_{0}=\beta_{\mathrm{yc}}(c^{\prime},\max\{d,d^{\prime}\},m){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}>1}$. ∎
@@ 4. Sard properties for infinite-dimensional maps
Given a relatively compact $S\subset\mathbb{R}^{m}$, for any $\varepsilon>0$ we denote with $M(\varepsilon,S)$ the $\varepsilon$-entropy of $S$, which is the minimum number of closed balls of radius $\varepsilon$ that we need to cover $S$.
@@
[…]
@@ 4.1. Sard-type theorem for well approximated maps
The purpose of this section is to prove the following theorem, which is our main Sard-type result (Theorem D in the Introduction).
@@
Let $d,m\in\mathbb{N}$. There exists a constant $\beta_{0}=\beta_{0}(d,m)>1$ such that the following holds. Let $H$ be a Hilbert space, and let $f:H\to\mathbb{R}^{m}$ be a $\mathcal{C}^{1}$ map such that its differential $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous. Let $K\subset H$ be a bounded set with this approximation property: there exist a sequence $E_{n}\subset H$ of linear subspaces, $\dim(E_{n})=n$, and polynomial maps $f_{n}:E_{n}\to\mathbb{R}^{m}$ with uniformly bounded degree:
@@
(121) $\sup_{x\in K}\bigg{(}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|+\|\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}\bigg{)}\leq cq^{-n}.$
@@
In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property on $K$:
(123) $\mu\bigg{(}f(\operatorname{Crit}(f)\cap K)\bigg{)}=0.$
@@
For the benefit of the reader we provide a general outline of the strategy of the proof. Our goal is to estimate the “size” of the set $f(\operatorname{Crit}_{\nu}(f)\cap K)$. More precisely, we bound the $\varepsilon$-entropy of $f(\operatorname{Crit}_{\nu}(f)\cap K)$ for any $\varepsilon>0$, and then pass to the limit $\varepsilon\to 0$.
We observe that the $\varepsilon$-entropy of $f(\operatorname{Crit}_{\nu}(f)\cap K)$ is controlled with the one of the set of almost-critical values $f(C^{\Lambda}(f)\cap K)$, see Definition 16. In turn, this is controlled by the $\varepsilon$-entropy of suitable almost-critical values of the approximating maps $f_{n}$, namely $f(C^{\Lambda_{\varepsilon}}(f_{n})\cap K)$, see (134).
To proceed, we use the relation between the $\varepsilon$-entropy of sets in $\mathbb{R}^{m}$ and their Vitushkin variations (Definition 14), provided by [YC04, Thm. 3.5].
Finally, we estimate the Vitushkin variations of $f_{n}(C^{\Lambda_{\varepsilon}}(f_{n})\cap K)$ applying Theorem G, exploiting the hypothesis that the maps $f_{n}$ are polynomial. This is the connection with quantitative semialgebraic geometry from Section 3. The crucial point of Theorem G is that the bound is quantitative with respect to parameters $\Lambda=\Lambda_{\varepsilon}$ and $n$. This allows to chose a large enough $n=n_{\varepsilon}$, see (148), to ensure convergence when passing to the limit $\varepsilon\to 0$.
@@
Step 1. We show that for any $\nu=0,\dots,m-1$ the set $\operatorname{Crit}_{\nu}(f)\cap K$ is contained in a suitable set of almost-critical points of $f$. By assumption, $K$ is weakly pre-compact. Since $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous, and by Lemma 17 the function $\sigma_{i}:\mathcal{L}(H,\mathbb{R}^{m})\to\mathbb{R}$ is continuous for every $i=1,\dots,m$. Therefore we have
(124) $\Sigma_{i}:=\sup_{K}\sigma_{i}(D_{x}f)<\infty.$
@@
(125) $\operatorname{Crit}_{\nu}(f)\cap K\subset C^{\Lambda}(f)\cap K.$
Step 2. We now relate critical points of $f$ to almost-critical points of the approximating polynomials $f_{n}$. Fix $\varepsilon>0$. Recall that, by Lemma 17, the singular values are $1$-Lipschitz with respect to the operator norm. Then for any sufficiently large $n\in\mathbb{N}$ such that $cq^{-n}\leq\varepsilon$ and for all $x\in K$ we have
@@
Hence, for all $x\in K\cap C^{\Lambda}(f)$ we have the following estimate for all $i=1,\dots,m$
(127) $\sigma_{i}\left(D_{\pi_{E_{n}}(x)}f_{n}\right)\leq\varepsilon+\sup_{y\in C^{\Lambda}(f)\cap K}\sigma_{i}\left(D_{y}f|_{E_{n}}\right)\leq\varepsilon+\sup_{y\in C^{\Lambda}(f)\cap K}\sigma_{i}\left(D_{y}f\right)\leq\varepsilon+\Lambda_{i},$
@@
(129) $\pi_{E_{n}}\left(C^{\Lambda}(f)\cap K\right)\subset C^{\Lambda^{\varepsilon}}(f_{n})\cap\pi_{E_{n}}(K)\subset C^{\Lambda^{\varepsilon}}(f_{n})\cap B_{E_{n}}(r),$
where $r>0$ is such that $K\subseteq B_{H}(r)$, and thus $B_{E_{n}}(r)=B_{H}(r)\cap E_{n}$ denotes the ball in $E_{n}$ defined by the restriction of the Hilbert norm.
Step 3. Now we use the inclusion (129) to estimate the entropy dimension of $f(\operatorname{Crit}_{\nu}(f)\cap K)$ with the one of almost-critical values of $f_{n}$. By (121), for sufficiently large $n$, we have
(130) $\sup_{x\in K}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|\leq\varepsilon.$
@@
(131) $M(2\varepsilon,f(\operatorname{Crit}_{\nu}(f)\cap K)$ $\leq M(2\varepsilon,f(C^{\Lambda}(f)\cap K))$ by (125)
(132) $\leq M\left(2\varepsilon,\mathcal{U}_{\varepsilon}\left(f_{n}\circ\pi_{E_{n}}\left(C^{\Lambda}(f)\cap K\right)\right)\right)$ by (130)
(133) $\leq M\left(\varepsilon,f_{n}\circ\pi_{E_{n}}\left(C^{\Lambda}(f)\cap K\right)\right)$ by definition of $\mathcal{U}_{\varepsilon}$
@@
The rest of the proof consists in estimating (134). We will use the results from Section 3.4, for this reason in the next step we discuss how to bring the problem to the Euclidean space $\mathbb{R}^{n}$.
@@
where $C(m)>0$ is a constant that depends only on $m$. We now apply the results from Section 3.4 in order to estimate the right-hand side of the last inequality.
Step 5. Take $n$ sufficiently large so that it also satisfies $n\geq m$. Then, by Theorem G, for every $i=0,\dots,m$ we have
@@
Step 6. We now show how to use (147) in order to get the estimate on $\dim_{e}(f(\operatorname{Crit}_{\nu}(f)\cap K))$. For $\varepsilon>0$ we choose $n=n_{\varepsilon}$ where
@@
(149) $\dim_{e}f(\operatorname{Crit}_{\nu}(f)\cap K)\leq\limsup_{\varepsilon\to 0}\frac{\ln M(\varepsilon,\tilde{f}_{n_{\varepsilon}}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n_{\varepsilon}})\cap B_{\mathbb{R}^{{n_{\varepsilon}}}}(r)))}{\ln\frac{1}{2\varepsilon}}.$
@@
(153) $\dim_{e}f(\operatorname{Crit}_{\nu}(f)\cap K)\leq\nu+\frac{\ln\beta_{0}}{\ln q},$

### v3

@@ 3.3. Variations and their behaviour under polynomial maps
Exactly as in [ComteYomdin, Thm. 7.2], we cannot apply directly the area formula to $A$, since the dimension of $A$ can be larger than $i$. In [ComteYomdin, Thm. 7.2] this problem is solved using [ComteYomdin, Ex. 4.11]. However, the dependence on $n$ of the constants appearing there is not explicit. It could in be principle be made so, using the Tarski-Seidenberg theorem, but the dependence obtained in this way would be doubly exponential in $n$, which is not enough for our purposes. We overcome the obstacle using Theorem˜14.
For sufficiently small $\varepsilon>0$ let $C_{\varepsilon}$ the semialgebraic set obtained by applying Theorem˜14 to the set $A$ and the polynomial map $F=p_{Z}:\mathbb{R}^{n}\to\mathbb{R}^{i}$. We then proceed using the area formula to estimate $\mathcal{H}^{i}(\pi_{Z}\circ p(C_{\varepsilon}))$. (We will see at the end of the proof how $\mathcal{H}^{i}(\pi_{Z}\circ p(C_{\varepsilon}))\to\mathcal{H}^{i}(\pi_{Z}\circ p(A))$ as $\varepsilon\to 0$). By Item˜i of Theorem˜14, $\dim(C_{\varepsilon})\leq i$. We can assume without loss of generality that $C_{\varepsilon}\subset\mathbb{R}^{n}$ is a smooth embedded submanifold with dimension equal to $i$, since the other strata give zero contribution to $\mathcal{H}^{i}$. Denoting by $\bar{p}_{Z}:C_{\varepsilon}\to\mathbb{R}^{i}$ the restriction, we obtain
@@
(86) $\leq\int_{C_{\varepsilon}}\sigma_{1}(D_{x}\bar{p}_{Z})\cdots\sigma_{i}(D_{x}\bar{p}_{Z})\mathcal{H}^{i}(dx).$ by definition of sing. values
Note that, denoting by $j:C_{\varepsilon}\hookrightarrow\mathbb{R}^{n}$ the inclusion, we have
@@
(90) $\leq\sup_{x\in\mathcal{U}_{\varepsilon}(A)}\sigma_{1}(D_{x}p)\cdots\sigma_{i}(D_{x}p)\mathcal{H}^{i}(C_{\varepsilon})$ by Item˜ii of Theorem˜14
(91) $\leq\left(\overline{\sigma}_{i}+\eta(\varepsilon)\right)\mathcal{H}^{i}(C_{\varepsilon}),$ by continuity of sing. values
where $\overline{\sigma}_{i}$ is defined in (77), and $\eta(\varepsilon)\to 0$ as $\varepsilon\to 0$. Since $C_{\varepsilon}$ is a bounded semialgebraic set of dimension $i$, then $\mathcal{H}^{i}(C_{\varepsilon})=V_{i}(C_{\varepsilon})$ by Lemma˜16. Using Items˜iv and ii of Theorem˜14, we obtain
@@
(93) $\leq c(n,i)\beta_{\mathrm{dc}}^{n-i}\int_{G_{i}(\mathbb{R}^{n})}\left(\int_{Z}\mathbbm{1}_{\pi_{Z}(C_{\varepsilon})}(x)\,\mathcal{H}^{i}(dx)\right)\gamma_{i,n}\left(dZ\right)$
(94) $\leq c(n,i)\beta_{\mathrm{dc}}^{n-i}\int_{G_{i}(\mathbb{R}^{n})}\left(\int_{Z}\mathbbm{1}_{\pi_{Z}(\mathcal{U}_{\varepsilon}(A))}(x)\,\mathcal{H}^{i}(dx)\right)\gamma_{i,n}\left(dZ\right)$
@@
where $\beta_{\mathrm{dc}}=\beta_{\mathrm{dc}}(c,d,i)>1$ is the constant from Theorem˜14, which (up to taking the maximum for $i\leq m$) depends only on $c,d,m$. Summing up, we have proved that for all $i=1,\dots,m$ it holds
@@
In order to take the limit in (97), we recall two properties of semialgebraic sets. The first one is the continuity of the Lebesgue measure in the Hausdorff topology in the semialgebraic category (see [ComteYomdin, Thm. 5.10]): if $\{S_{\varepsilon}\}_{\varepsilon\geq 0}\subset\mathbb{R}^{i}$ is a one-parameter family of bounded semialgebraic sets then
@@
The second property is the behaviour of variations for $\varepsilon$-neighbourhoods (see [ComteYomdin, Thm. 5.11]): if $A\subset\mathbb{R}^{n}$ is a bounded semialgebraic set, then for all $i=1,\dots,n$ it holds
@@
Therefore, using (98) (and recalling that by Item˜iii of Theorem˜14 it holds $p_{Z}(C_{\varepsilon})\to p_{Z}(A)$ in the Hausdorff topology) and (99) we can pass to the limit in (97) and obtain
@@
To conclude, we estimate $V_{i}(A)$. Similarly as in (83), since $A\subset B_{\mathbb{R}^{n}}(r)$ we have
@@ 3.4. Variations of almost-critical values of polynomial maps
In this section we recall the notion of almost-critical points $C^{\Lambda}(p)$ of a polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$. The set $p(C^{\Lambda}(p))$ is the set of its almost-critical values. The main result of this section is Theorem˜26, which provides a version of [ComteYomdin, Cor. 7.4] where the dependence of the parameters with respect to $n$ is explicit. This result is based on Theorem˜19 and on the specific semialgebraic structure of almost-critical points of polynomial maps, that we prove in Proposition˜25.
@@
where $E_{i,b_{i}+1}=\{(x,y)\in\mathbb{R}^{s}\times\mathbb{R}\mid y-t\leq 0\}$. In particular, this set admits a diagram that does not depend on $t$, since the degree of the polynomial yty-t is equal to $1$. As explained in Section˜3.1.3, the projection of a semialgebraic set is semialgebraic and its diagram is controlled explicitly by the diagram of the original set only. In particular, the diagram of the projection (107) admits a description with a diagram $(s,I,J)$, that does not depend on $t$. ∎
@@
The functions $\lambda_{i}:{\operatorname{Sym}}_{m}^{+}\to\mathbb{R}$ are continuous and semialgebraic.
@@
By Lemma˜24, for every $k=1,\ldots,m$, the function $\lambda_{k}:{\operatorname{Sym}}_{m}^{+}\to\mathbb{R}$ is semialgebraic, hence from Lemma˜23 applied to $A=\operatorname{Sym}_{m}^{+}$ and the semialgebraic functions $\lambda_{k}:A\to\mathbb{R}$, we deduce the following fact: for any $t\in\mathbb{R}$ we can write
@@
Now we state and prove the main result of this section (Theorem˜G in the Introduction).
@@
By Proposition˜25 the set $A=C^{\Lambda}(p)\cap B_{\mathbb{R}^{n}}(r)$ is a closed and bounded semialgebraic set whose diagram is $D(A)=(n,c^{\prime},d^{\prime})$ where $c^{\prime}=c^{\prime}(d,m)$, $d^{\prime}=d^{\prime}(d,m)$ (which in particular do not depend on $n$, $\Lambda$). Furthermore, for the polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ it holds, by construction
@@
We can then apply Theorem˜19, yielding (115) with $\beta_{0}=\beta_{\mathrm{yc}}(c^{\prime},\max\{d,d^{\prime}\},m)>1$. ∎
@@ 4. Sard properties for infinite-dimensional maps
Given a relatively compact $S\subset\mathbb{R}^{m}$, for any $\varepsilon>0$ we denote by $M(\varepsilon,S)$ the $\varepsilon$-entropy of $S$, which is the minimum number of closed balls of radius $\varepsilon$ that we need to cover $S$.
@@
the operator norm. With some abuse of notation the same symbol $\|L\|_{\operatorname{op}}$ is used for different $X$ and $Y$, but there should be no confusion as the domain and codomain of $L$ are clear from the context. We also use without risk of confusion the symbol $\|\cdot\|$ to denote the usual Euclidean norm of $\mathbb{R}^{m}$.
@@ 4.1. Sard-type theorem for well approximated maps
The purpose of this section is to prove the following theorem, which is our main Sard-type result (Theorem˜D in the Introduction).
@@
Let $d,m\in\mathbb{N}$. There exists a constant $\beta_{0}=\beta_{0}(d,m)>1$ such that the following holds. Let $H$ be a Hilbert space, and let $f:H\to\mathbb{R}^{m}$ be a $\mathcal{C}^{1}$ map such that its differential $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous. Let $B\subset H$ be a bounded set with this approximation property: there exist a sequence $E_{n}\subset H$ of linear subspaces, $\dim(E_{n})=n$, and polynomial maps $f_{n}:E_{n}\to\mathbb{R}^{m}$ with uniformly bounded degree:
@@
(123) $\sup_{x\in B}\bigg(\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|+\|\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}\bigg)\leq cq^{-n}.$
@@
(124) $\dim_{e}\bigg(f\big(\operatorname{Crit}_{\nu}(f)\cap B\big)\bigg)\leq\nu+\frac{\ln\beta_{0}}{\ln q}.$
In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property on $B$:
(125) $\mu\bigg(f(\operatorname{Crit}(f)\cap B)\bigg)=0.$
@@
For the benefit of the reader we provide a general outline of the strategy of the proof. Our goal is to estimate the “size” of the set $f(\operatorname{Crit}_{\nu}(f)\cap B)$. More precisely, we bound the $\varepsilon$-entropy of $f(\operatorname{Crit}_{\nu}(f)\cap B)$ for any $\varepsilon>0$, and then pass to the limit $\varepsilon\to 0$.
We observe that the $\varepsilon$-entropy of $f(\operatorname{Crit}_{\nu}(f)\cap B)$ is controlled with the one of the set of almost-critical values $f(C^{\Lambda}(f)\cap B)$, see Definition˜17. In turn, this is controlled by the $\varepsilon$-entropy of suitable almost-critical values of the approximating maps $f_{n}$, namely $f(C^{\Lambda_{\varepsilon}}(f_{n})\cap B)$, see (136).
To proceed, we use the relation between the $\varepsilon$-entropy of sets in $\mathbb{R}^{m}$ and their Vitushkin variations (Definition˜15), provided by [ComteYomdin, Thm. 3.5].
Finally, we estimate the Vitushkin variations of $f_{n}(C^{\Lambda_{\varepsilon}}(f_{n})\cap B)$ applying Theorem˜G, exploiting the hypothesis that the maps $f_{n}$ are polynomial. This is the connection with quantitative semialgebraic geometry from Section˜3. The crucial point of Theorem˜G is that the bound is quantitative with respect to parameters $\Lambda=\Lambda_{\varepsilon}$ and $n$. This allows to chose a large enough $n=n_{\varepsilon}$, see (150), to ensure convergence when passing to the limit $\varepsilon\to 0$.
@@
Step 1. We show that for any $\nu=0,\dots,m-1$ the set $\operatorname{Crit}_{\nu}(f)\cap B$ is contained in a suitable set of almost-critical points of $f$. By assumption, $B$ is weakly pre-compact. Since $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous, and by Lemma˜18 the function $\sigma_{i}:\mathcal{L}(H,\mathbb{R}^{m})\to\mathbb{R}$ is continuous for every $i=1,\dots,m$. Therefore we have
(126) $\Sigma_{i}:=\sup_{B}\sigma_{i}(D_{x}f)<\infty.$
@@
(127) $\operatorname{Crit}_{\nu}(f)\cap B\subset C^{\Lambda}(f)\cap B.$
Step 2. We now relate critical points of $f$ to almost-critical points of the approximating polynomials $f_{n}$. Fix $\varepsilon>0$. Recall that, by Lemma˜18, the singular values are $1$-Lipschitz with respect to the operator norm. Then for any sufficiently large $n\in\mathbb{N}$ such that $cq^{-n}\leq\varepsilon$ and for all $x\in B$ we have
@@
Hence, for all $x\in B\cap C^{\Lambda}(f)$ we have the following estimate for all $i=1,\dots,m$
(129) $\sigma_{i}\left(D_{\pi_{E_{n}}(x)}f_{n}\right)\leq\varepsilon+\sup_{y\in C^{\Lambda}(f)\cap B}\sigma_{i}\left(D_{y}f|_{E_{n}}\right)\leq\varepsilon+\sup_{y\in C^{\Lambda}(f)\cap B}\sigma_{i}\left(D_{y}f\right)\leq\varepsilon+\Lambda_{i},$
@@
(131) $\pi_{E_{n}}\left(C^{\Lambda}(f)\cap B\right)\subset C^{\Lambda^{\varepsilon}}(f_{n})\cap\pi_{E_{n}}(B)\subset C^{\Lambda^{\varepsilon}}(f_{n})\cap B_{E_{n}}(r),$
where $r>0$ is such that $B\subseteq B_{H}(r)$, and thus $B_{E_{n}}(r)=B_{H}(r)\cap E_{n}$ denotes the ball in $E_{n}$ defined by the restriction of the Hilbert norm.
Step 3. We use the inclusion (131) to estimate the entropy dimension of $f(\operatorname{Crit}_{\nu}(f)\cap B)$ with the one of almost-critical values of $f_{n}$. By (123), for sufficiently large $n$, we have
(132) $\sup_{x\in B}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|\leq\varepsilon.$
@@
(133) $M(2\varepsilon,f(\operatorname{Crit}_{\nu}(f)\cap B)$ $\leq M(2\varepsilon,f(C^{\Lambda}(f)\cap B))$ by (127)
(134) $\leq M\left(2\varepsilon,\mathcal{U}_{\varepsilon}\left(f_{n}\circ\pi_{E_{n}}\left(C^{\Lambda}(f)\cap B\right)\right)\right)$ by (132)
(135) $\leq M\left(\varepsilon,f_{n}\circ\pi_{E_{n}}\left(C^{\Lambda}(f)\cap B\right)\right)$ by definition of $\mathcal{U}_{\varepsilon}$
@@
The rest of the proof consists in estimating (136). We will use the results from Section˜3.4, for this reason in the next step we discuss how to bring the problem to the Euclidean space $\mathbb{R}^{n}$.
@@
where $C(m)>0$ is a constant that depends only on $m$. We now apply the results from Section˜3.4 in order to estimate the right-hand side of the last inequality.
Step 5. Take $n$ sufficiently large so that it also satisfies $n\geq m$. Then, by Theorem˜G, for every $i=0,\dots,m$ we have
@@
Step 6. We show how (149) yields the estimate on $\dim_{e}(f(\operatorname{Crit}_{\nu}(f)\cap B))$. For $\varepsilon>0$ we choose $n=n_{\varepsilon}$ where
@@
(151) $\dim_{e}f(\operatorname{Crit}_{\nu}(f)\cap B)\leq\limsup_{\varepsilon\to 0}\frac{\ln M(\varepsilon,\tilde{f}_{n_{\varepsilon}}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n_{\varepsilon}})\cap B_{\mathbb{R}^{{n_{\varepsilon}}}}(r)))}{\ln\frac{1}{2\varepsilon}}.$
@@
(155) $\dim_{e}f(\operatorname{Crit}_{\nu}(f)\cap B)\leq\nu+\frac{\ln\beta_{0}}{\ln q},$